OpenStudy (anonymous):

write radical expression for -64^3/2

OpenStudy (anonymous):

The denominator of your exponent is the number outside the radical. The numerator stays with the -64. \[\sqrt[2]{-64^{3}}\]

OpenStudy (anonymous):

You can simplify by raising -64 for the 3rd power, and dropping the 2 outside the radical, as it is understood. Below are equivalent expressions. \[\sqrt{-262144}\] and \[\sqrt{-64^{3}}\]